# Can Robots “Taste” Grapes? Estimating SSC with Simple RGB Sensors

### v1

@@ 3.1.2 Color Constancy effect on Human Evaluation of Color
One of the early questions that arose from data collection was to what extent the camera image processing pipeline could influence the possibility of estimating the correct color stage, and therefore influence the °Brix estimate. Smartphone cameras, like most camera devices, do not process the raw signal from their CCD/CMOS sensors for radiation accuracy, but more for aesthetic purposes. This means that color accuracy is often lost in the nonlinear pipeline that transforms raw images into JPGs. A second consideration was related to the difference between the human vision system and a camera device. We know that human experts and workers alike are able to guess ripeness from appearance, but to what extent some of the peculiarities of the human vision system influence this guess (e.g. consider the brightness constancy effect) is unknown. While color perception is a complex topic, we focused our analysis on a specific question: whether human experts assign significantly different color labels when viewing white-balanced versus standard images. We selected 50 images from a single field trip that had paired raw and JPG image formats. The raw image was processed only with debayering and white balancing using the grey card. (Figure 7 shows some examples).
Figure 7: Some examples of pictures of the same table grape bunches were presented to agronomists on different occasions to test whether human experts consistently evaluated bunch ripeness, regardless of color correction, as described in Section 3.1.2
We presented all 100 images in random sequence to a group of agronomist experts, without revealing that each bunch appeared twice. The experts labeled the color quality of each image using the color scale described in Section 3, without performing any other manipulation or consideration. The null hypothesis is that there is no difference in labeling standard and white-balanced images apart from some generic labeling noise. If we define the mean group labeling error as:
$\mu_{i}=\mathbf{E}\left[L_{GT_{i}}-L_{exp_{i}}\right]$ (7)
where $i\in[0,1]$ is the sample group, $L_{GT_{i}}$ is the hypothetical ground truth color label for a random sample from group $i$, and $L_{exp_{i}}$ is the label that an expert gives for the same random sample. Furthermore, we denote with $Y_{i}$ the sample mean of the difference between the true label and the estimated one. Then, we can state that $Y_{1}-Y_{0}$ is an unbiased normally distributed estimator of $\mu_{1}-\mu_{0}$, and thus we can formulate the following null hypothesis:
$\mathcal{H}_{0}:\mu_{1}-\mu_{0}=0$ (8)
which we can easily test for significance with a z-test.
Analysis of the results revealed only a small difference between expert labels for the two groups, with a mean difference of $0.1$ and a standard deviation of $0.50$. The paired test for the sample batch gives a z-statistic of 1.4 (p=0.17), which does not provide sufficient evidence to reject $\mathcal{H}_{0}$. While a larger sample size might yield more definitive results, we chose instead to focus on the estimation algorithms, which are the main subject of our study. In particular, we included the comparison between white-balanced images and standard images into the cross-validation of the histogram-based features to gain deeper insights.
3.2 Histogram based color and °Brix estimation
This Section first outlines the experimental setup used to evaluate histogram features and subsequently presents the results.
3.2.1 Description
Given the exploratory experiments described in the previous Sections 3.1.1 and 3.1.2, we designed the histogram features as described in Section 2.2 and use them to train regression and classification models to predict SSC and color level, respectively.
The following experiments were conducted for both color estimators (multi-class classification) and SSC estimators (°Brix regression). Given the higher difficulty of SSC estimation compared to exterior color evaluation, the following description focuses on SSC estimation, while the tables and figures summarize the results for both SSC and color estimation.
The main question addressed in these experiments is the overall performance of the histogram-based algorithms and the influence of hyper-parameters. A number of grid hyper-parameter searches using Ridge Regression were conducted on a limited dataset, specifically the $T1_{STD}$ and $T1_{WB}$ splits, to mitigate distribution shifts due to different camera sensors and seasons during the initial analysis. These splits contain 112 samples each, divided into 89 train images and 23 validation images. The training images were augmented with horizontal flips, resulting in a total of 178 training images. As previously mentioned, both standard JPG and white-balanced images were evaluated to assess the influence of post-processing on the feature-based algorithms.
To set a baseline for this approach, we considered the intrinsic variability training and test data. The $T1$ datasets have a mean of 17.19 °Brix, a standard deviation of 2.43 °Brix, and a mean absolute deviation (MAD) of 1.99 °Brix. Appendix A provides more details on the statistical properties of the datasets. The naive predictor, which always predicts the mean °Brix values of the training set, resulting in a mean absolute error (MAE) equal to the MAD, serves as the lower bound for estimator performance.
@@ 3.2.1 Description
Figure 8: a) Correlation between validation error and hyper-parameters for the top 10% performing runs. This plot highlights the direction (positive or negative) of the correlation. b) Test error for models with hyper-parameters identical to the best model shown in Table 3, but with varying $alpha$ values and color correction. The consistent lower error for models trained on standard images compared to color-corrected images gives a clear answer to the question raised in Section 3.1.2.
Test MAE Val MAE Train MAE $alpha$ $n\_bin\_x$ $n\_bin\_y$ $cross$ $color\_space$ $feature\_len$ $log(alpha)\times feature\_length$
1.459898 2.098344 1.583124 9.000000 16 8 none hsv 128 122.143041
1.509351 2.129067 1.548176 9.000000 8 16 none hsv 128 122.143041
1.522208 2.141015 1.614265 3.000000 8 8 none hsv 64 30.535760
1.509816 2.144949 1.178387 9.000000 8 32 none hsv 256 244.286082
1.507421 2.164503 1.619756 1.000000 16 8 thin lab 80 0.000000
1.495739 2.164640 1.254232 3.000000 16 8 none hsv 128 61.071521
1.517718 2.181812 1.431866 3.000000 16 8 fat hsv 96 45.803640
1.523156 2.193494 1.719893 0.330000 8 8 thin lab 40 -19.259442
1.431197 2.200048 1.189874 3.000000 8 16 none hsv 128 61.071521
1.491955 2.327332 1.225366 1.000000 16 8 thin hsv 80 0.000000
Table 3: The top 10 histogram-based models, ranked by validation mean absolute error (Val MAE), along with their corresponding hyper-parameters values. Some trends discussed in Section 3.2, such as the importance of alternative color spaces like HSV, are clearly evident.
3.2.2 Discussion
A grid search on the aforementioned hyper-parameter ranges, requiring training 3888 different models, was performed using the Weights & Biases experiment tracking framework [2] to compare and analyze the results. Further details of the hyper-parameter search are provided in Appendix B, while this section focuses on the key findings. The correlation coefficient of each feature with the validation error has been computed and is shown in Fig. 8(a). In addition, Table 3 shows the 10 best runs for the $low\_res$ images.
As expected, given the limited number of samples in the $T1$ datasets, the regularization parameters and those influencing feature vector length had the greatest impact on the final results. Furthermore, image resolution proved irrelevant, suggesting the safe use of lower-resolution images for computational efficiency in embedded applications. This is consistent with the feature construction process, where binning and averaging naturally remove the complexities of high-level details. Finally, and most notably, both Fig.8(a) and Table 3, indicate that using an alternative color space, such as HSV, helps in finding better coefficients for the Ridge Regression.
Focusing solely on the 1296 lower-resolution runs, the best-performing models have validation MAE of $1.45$, with different combinations of hyper-parameters that have in common a low feature count relative to the regularization value. Table 3 summarizes some of these top-performing runs, all exhibiting MAE lower than the baseline brixmad of 1.99. The best run reaches a $\mathbf{27\%}$ improvement over the baseline, a notable result considering the limited training data. Figure 8(b) displays the validation errors for different $alpha$ values (on a logarithmic scale), for both color-corrected and standard images. The most interesting feature of the plot, which is a common trend in all similarly produced plots, is the consistently inferior performance of the models trained on color-corrected images compared to their standard counterparts with identical hyper-parameters.
3.3 Multi-task Learning DNN
The SSC and color distribution can be viewed as conditioned on the environmental and agronomic parameters of the field we are considering. While both the color and the SSC depend on the internal chemical state of the berries, this state itself is influenced by environmental factors such as growing conditions, training practices, and nutrient availability, The overall appearance of the grape bunches is also affected by these factors, which we assume are homogeneous across the vineyard.
The architecture of the multi-task network, illustrated in Fig. 5, reflects this reasoning. The encoder component captures common features in appearance variations that may correlate with the internal chemical state of the grapes. These shared features are then leveraged by two separate prediction heads to estimate the color class and SSC in parallel. As before, in the following we first give a description of the experiment and then discuss the results.
3.3.1 Description
The $T2$ dataset (Table 1) is used to evaluate the covariate shift due to the variability in treatments between consecutive seasons while keeping the sensor type fixed. Since phone images were captured in both raw and JPG formats, the $T2$ splits include standard ($T2_{STD}$) and white-balanced ($T2_{WB}$) versions. This allows for a direct comparison of performances between standard JPG and color-corrected images, under identical hyper-parameters and for the same grape bunch instances. In both cases, data from the 2021 season serves as the training and validation sets, while the data from the 2022 season are used as the test set. The $T2_{ALL}$ split consists only of color-corrected images since it includes reflex camera images that have been captured in raw format and subsequently color-corrected using the gray card. This split helps assess to some degree the effect of using two different camera devices to capture the same instances with the same perspective. The $T3$ splits combine images from all devices across both seasons, including both color-corrected and standard versions.
Appendix C and Table 10 summarize the hyper-parameter ranges considered for the experimental evaluation. We performed $4912$ runs on a DGX server with four $A100$ GPUs. Each run was trained for $200$ epochs with hyper-band early stopping. For all the experiments we used a ResNet18 backbone [7], exploring both the effect of pretraining on ImageNet [4] and training from scratch. Hyper-parameter relevance was assessed using the BCH metric described in Section 2.3.1. While the complete details on the analysis of the importance of each hyper-parameter are described in Appendix C, in the following Section we discuss the most relevant findings.
Figure 9: Multi-task network hyper-parameter correlation - These histograms shows the correlation between the validation error and the hyper-parameters computed on the 25% best performing runs. a) $T2$ dataset: it can be noted that the main influence comes from the classical parameters that impact any DNN training (learning rate, batch size, regularization parameters), though image size and color space remain relevant. More insights on this can be extracted from Table 4. b) $T3$ dataset: similar trends are observed as in (a), with classical DNN training parameters having the most significant impact. The increased importance of blur parameters suggests the potential benefits of data augmentation strategies in mitigating the effects of covariate shifts due to variations in camera devices.
3.3.2 Discussion
We begin our analysis with the $1002$ runs performed on the $T2$ splits. Since we sampled a wide range of models via random search, we concentrate on the best $25\%$ of the total runs to discuss the influence of hyper-parameters. Figure 9(a) presents the correlation between the hyper-parameters and the BCH metric, while Table4 lists the top $10$ performing runs. The multi-task network effectively learns good models for both SSC and color estimation. Apart from the typical learning parameters, the ones that show more influence on the performance results are image size (particularly width) and color space (rarely RGB among top performers). However, no single hyper-parameter shows high correlation values, indicating a complex relationship linking the parameters to the error. Another consideration is that $\alpha$ is typically ten times larger than $\beta$, giving a rough estimate of the relative weight of the corresponding losses for the final results. Moreover, the benefit of the additional augmentations appears less substantial than initially anticipated. It is important to note that the lowest MAE for SSC and the highest accuracy for color estimation are not achieved at the same epoch. This is unsurprising, as perfectly correlated feature learning for distinct tasks is unlikely, even with strong underlying correlations. Therefore, when using a single feature extractor, a trade-off must be made between optimizing for SSC or color estimation. If computational resources permit, employing two different feature extractors (i.e., different checkpoints) allows for independent optimization and potentially better performance on both tasks.
Val BCH Val SSC MAE Val Acc $alpha$ $beta$ $im\_width$ $im\_height$ $momentum$ $learning\_rate$ $batch\_size$ $blur\_\sigma\_min$ $blur\_\sigma\_max$ $pretrained$ $one\_cycle$ $HSV$ $Lab$ $augment\_extra$
4.076824 1.555511 0.673913 500 50 256 256 0.990000 0.000001 64 0.100000 2 False False False True True
4.188802 1.832541 0.663043 1000 1000 256 64 0.950000 0.000001 32 0.500000 4 False True False True False
4.265176 1.653810 0.554348 100 10 64 256 0.950000 0.000050 64 0.500000 2 True False True False False
4.312880 1.732056 0.619565 500 100 256 64 0.800000 0.000050 128 0.100000 2 True True True False True
4.346299 1.579317 0.630435 1000 100 256 128 0.800000 0.000050 128 0.010000 4 False True True False False
4.412035 2.160111 0.532609 500 100 64 128 0.990000 0.000050 128 0.010000 2 False True True False False
4.453714 1.795929 0.684783 1000 500 256 256 0.950000 0.000010 128 0.010000 2 False False False True True
4.496901 1.891276 0.673913 100 50 256 128 0.900000 0.000050 128 0.500000 2 True True True False True
4.536293 1.629490 0.521739 100 100 256 128 0.900000 0.000050 64 0.500000 2 False False False False False
4.598450 1.620318 0.695652 10 1 256 256 0.990000 0.000050 128 0.100000 4 False True False True True
Table 4: The best ten runs on $T2$ split.
Similar trends emerge from the experiments on the $T3$ splits. Figure 9(b) shows the correlations of the hyper-parameters with the error, and Table 13 displays the best-performing runs. Again, the $\alpha$ to $\beta$ ratio is confirmed, along with the general preference for having higher image width values. The $T3$ test set, captured with a Realsense d435i, introduces a covariate shift due to its lower resolution and different white balance settings compared to the phone and reflex cameras used for training. This difference might explain the increased relevance of blur-related parameters in the augmentation.
Following the analysis conducted for histogram features in Section 3.2, we compare the best $T3$ run with runs using the same hyper-parameters but trained on white-balanced images. Unlike the histogram feature case with a single regularization parameter, the multi-task model utilizes two ($\alpha$ and $\beta$). Given what was observed from the experiments, that the ratio of the $\alpha$ and $\beta$ values is more important than their absolute value, we use this as a parameter to compare the impact of using color-corrected images versus using the standard JPG ones. Similarly to the analysis presented in Section 3.2, we select the 150 best-performing runs and the ones with matching hyper-parameters (except for $\log(\alpha/\beta)$) and use them to extract a general trend shown in the scatter plot in Figure 10. The regression lines for both standard and white-balanced image sets exhibit nearly identical slopes ($-0.212$ and $-0.216$, respectively), and the difference between the lines is small, with largely overlapping confidence intervals. This suggests that the multi-task model effectively extracts relevant features regardless of color correction, with performance oscillations likely stemming from training data variability and other hyper-parameters.
To conclude, the best-performing runs for SSC estimation achieve MAE values as low as 1.55 (1.57) for the model trained on the $T2_{STD}$ ($T2_{WB})$ split, and 1.11 (1.10) for the one trained on the $T3_{STD}$ ($T3_{WB})$ split. Considering that the MAD values for the $T2$ and $T3$ test splits are 1.71 and 1.36, respectively, the performance gain of the multi-task models is $9.3\%$ ($8.1\%$) for the $T2$ case and $18.4\%$ ($19.1\%$) for $T3$ case. These results confirm the negligible impact of color correction on the multi-task model performance, as observed in the regression analysis. The superior performance of the white-balanced model on the $T3$ dataset can be attributed to the automatic white balancing performed by the d435i camera used for collecting the $T3$ test set.
@@ 3.3.2 Discussion
Figure 10: Linear regression analysis of runs with hyper-parameters similar to the top 10 runs on the $T3_{STD}$ and $T3_{WB}$ datasets. The shaded area represents the $95\%$ confidence interval. While the models trained on color-corrected samples present slightly higher values on average, the smaller difference compared to Fig. 8(b) suggests no practical advantage in using standard images over the white-balanced ones in this context.
@@ 4 Conclusions
This work addressed the problem of estimating soluble solid content (SSC) in grapes using RGB data acquired in the field. While numerous studies explore laboratory equipment and hyperspectral data for SSC estimation through visual sensors, this study is among the few to focus exclusively on the practical scenario of RGB data collected directly in field conditions. The complexity of this task arises from covariate shifts inherent to the problem, including seasonal variations and differences in imaging devices.
@@
Two algorithmic approaches were explored: a light-weight method using classical handcrafted features, suitable for devices with limited computational power, and a multi-task neural network architecture that, while more computationally demanding, effectively leverages correlations between visual appearance (color and shades) and SSC. Both methods were experimentally validated using field-collected data from the Pizzutello Nero grape variety. Results indicate that the proposed techniques achieve reliable SSC estimation with a reasonable margin compared to the average human estimator. Furthermore, experiments demonstrated some robustness to covariate shifts, including seasonal variations and device differences, positioning the proposed solutions as viable prototypes for low-cost, on-field applications where frequent data collection and labeling are challenging.
We release all the data, code, trained models, and training scripts used in this study. Future work will focus on refining and aligning these low-cost estimators and benchmarking them against hyperspectral approaches.

### v2

@@ 3.1.2 Color Constancy effect on Human Evaluation of Color
As part of our exploratory analysis, we investigated whether the camera’s automatic post-processing (standard JPG) versus a color-corrected image (white-balanced RAW) would significantly affect a human expert’s color evaluation. To test this, we conducted a study where agronomists labeled 50 pairs of images (one standard, one white-balanced) in a randomized sequence. A statistical analysis (z-test) of the assigned labels revealed no significant difference between the two groups (p = 0.17). This null finding confirmed that for human experts, the two image types were perceptually similar and justified our subsequent approach of testing our computer vision models on both standard and white-balanced images to evaluate their robustness.
3.2 Histogram-based Model Evaluation
To establish a performance baseline for a computationally inexpensive approach, we evaluated the histogram-based feature model using Ridge Regression. The experiments were conducted on the Simple dataset splits, which contain 112 samples. For this analysis, the data was divided into a training set (89 samples) and a final test set (23 samples). The training images were augmented with horizontal flips, resulting in a total of 178 training images. As previously mentioned, both standard JPG and white-balanced images were evaluated to assess the influence of post-processing on the feature-based algorithms. Model performance and hyper-parameters were then assessed using a 5-fold cross-validation procedure on the training set. The model’s performance was optimized by conducting a grid search over its key hyper-parameters. These included the number of horizontal and vertical histogram bins ($n_{bin_{x}}$, $n_{bin_{y}}$), the resulting feature vector length (feature_len), the choice of color space (RGB, HSV, or CIE-Lab), and the regularization strength ($\lambda$). The full list of parameters and their search ranges is detailed in A. The best-performing model achieved a validation MAE of $1.46$Brix. This result represents a notable $\mathbf{27\%}$ improvement over the statistical baseline for this dataset, which had a Mean Absolute Deviation (MAD) of $1.99$Brix.
The analysis of the hyper-parameter search revealed key insights into the model’s behavior. As shown in the feature correlation plot (Fig. 9(a)), the regularization strength ($\lambda$) and parameters affecting the feature_len had the most significant impact on performance. The analysis also highlighted the consistent benefit of using an alternative color space, with both HSV and CIE-Lab generally outperforming standard RGB.
Furthermore, we investigated the effect of using standard JPG images versus color-corrected, white-balanced (WB) images. Fig. 9(b) shows that models trained on the standard JPG images consistently achieved a lower validation error than their counterparts trained on white-balanced images with identical hyper-parameters. This suggests that for this feature-based approach, the camera’s default processing pipeline may contain beneficial information, and the extra step of color correction is not necessary.
@@ 3.2 Histogram-based Model Evaluation
Figure 9: a) Correlation between validation error (MAE) and hyper-parameters for the top 10% performing runs. This plot highlights the direction (positive or negative) of the correlation. Parameters controlling model complexity, such as feature_len and lambda (regularization strength), have the strongest correlation with performance. b) Test error for models with hyper-parameters identical to the best model, but with varying lambda values (plotted on a logarithmic scale) and color correction. This plot isolates the effect of regularization and demonstrates that, for this model, standard JPG images consistently yield better performance than their color-corrected counterparts.
3.3 Multi-task DNN Model Evaluation
[…]
An important observation from the multi-task training process is that the optimal performance for SSC estimation (regression) and color quality (classification) were not always reached at the same training epoch. This highlights an inherent trade-off when using a single feature extractor for multiple correlated but distinct tasks. This also implies that, for a specific application where the top performance of either SSC or color estimation is required, one could select different model checkpoints depending on whether the primary goal is maximizing SSC accuracy or color classification accuracy.
Finally, we analyzed the impact of color correction on the DNN model’s performance, similarly to what was done with the histogram features. To isolate this effect, we selected the best-performing models on the standard $\texttt{MixedSeason}_{\texttt{STD}}$ dataset and compared them against runs using identical hyper-parameters but trained on the white-balanced $\texttt{MixedSeason}_{\texttt{WB}}$ data. Fig. 10 plots the results of this comparison against the log-ratio of the regularization weights ($\alpha/\beta$). The resulting regression lines show that the performance difference between the two sets of images was minimal, with largely overlapping confidence intervals. This suggests that the deep learning model is robust to these color variations and can learn the relevant features effectively without requiring complex color calibration steps prior to deployment.
@@ 3.3 Multi-task DNN Model Evaluation
Figure 10: Linear regression analysis of runs with hyper-parameters similar to the top 20 runs on the $\texttt{MixedSeason}_{\texttt{STD}}$ and $\texttt{MixedSeason}_{\texttt{WB}}$ datasets. The shaded area represents the $95\%$ confidence interval. While the models trained on color-corrected samples present slightly higher values on average, the smaller difference compared to Fig. 9(b) suggests no practical advantage in using standard images over the white-balanced ones in this context.
4 Discussion
The results of our experiments demonstrate that low-cost, in-field RGB imaging can be a viable alternative to more complex and expensive systems for SSC estimation. In this section, we synthesize the performance of our models, compare them against the established baselines and the state-of-the-art, and discuss the practical implications and limitations of our work.
4.1 Summary of Performance and Comparison
To provide a holistic view of our findings, Table 7 consolidates the performance of our proposed models against the statistical and human baselines, as well as against other non-contact methods in the literature.
Table 7: Master comparison of model performance against baselines and state-of-the-art methods.
[…]
Our deep learning model clearly learned a meaningful pattern from the images, achieving a best MAE of 1.55 °Brix on the Season21ToSeason22Phone dataset and 1.05 °Brix on the challenging cross-device MixedSeason dataset. This significantly outperforms the naive statistical baselines of 1.71 and 1.36 °Brix MAD respectively for that same test set.
To better address how these errors translate to harvesting process performance and to benchmark against practical expertise, we also evaluated our DNN model on the binary harvesting task. As shown in Table 7, our model achieved a precision/recall of $\mathbf{0.91}/\mathbf{0.84}$ on Season21ToSeason22Phone split and of $\mathbf{0.62}/\mathbf{0.83}$ on the MixedSeason split. This is compared to the experts’ performance of $0.43/0.83$. Being able to keep a high recall, while showing a higher precision is particularly valuable in a real-world scenario, as it would reduce the costly error of harvesting unripe grapes. In addition, our experiments consistently showed that the F1 score computed from these Precision and Recall values is highly correlated to the BCH metric (in most runs the maxima of the F1 score correspond to the minima of the BCH metric) suggesting that the model’s high performance derives from the effective combination of features related to both color and SSC estimation.
When benchmarked against other visual estimation techniques, our low-cost RGB approach demonstrates highly competitive performance. While direct comparison is challenging due to different grape varieties and conditions, our best MAE of 1.05 °Brix is favorable when compared to reported RMSE values from in-field HSI systems, which range from 1.27 to 2.20 °Brix. It also represents a significant improvement over other in-field RGB methods that report errors closer to 4.63 °Brix RMSE, and approaches the accuracy of methods performed under controlled laboratory conditions.
4.2 Practical Implications and Trade-offs
Our work presents two distinct solutions tailored for different operational needs. The histogram-based model, while less accurate, offers significant advantages in computational efficiency and simplicity. Once trained, the estimation reduces to a dot product between the feature vector and the estimator weights. The most expensive operation in that case is the pixel-wise integral sum needed for the histogram computation, which has a computational complexity of O(N), where N is the number of pixels. However, as it was shown in the experiments, given the low frequency content of the grape image it is possible to work with very low resolution images and drastically reduce N. To summarize, the histogram features’ low memory and processing footprint make it suitable for real-time deployment on resource-constrained edge devices or as a first-pass filter on a robotic platform. In contrast, the DNN model provides superior accuracy and generalization, particularly across different devices and seasons, at a greater computational cost. This makes it ideal for applications where performance is paramount, such as a centralized quality control system or deployment on robots with dedicated GPU hardware.
4.3 Limitations and Future Work
We acknowledge several limitations that provide avenues for future research. This study was conducted on a single black grape variety, Pizzutello Nero. While this is common in literature (see Table 7), to extend it to a white variety some adjustments in the setup are required. In the case of black or red grape varieties, the color gradient can be perceived as a contrast change, while in with the white varieties it is a hue and saturation change. To detect variations of the latter kind, a color calibrated RGB imaging system is needed, or, in alternative, to have a calibration pattern in the scene. This is not impossible to achieve and it is an interesting future development for this work. Additionally, while our methods showed robustness to the uncontrolled lighting of the field, further experiments under more extreme weather conditions (e.g., direct bright sunlight vs. heavy clouds) could better quantify the operational domain of the system. Future work will focus on expanding the dataset to include these variations and on exploring multi-sensor fusion techniques to further improve robustness.
@@ 5 Conclusions
This work successfully demonstrated the feasibility of estimating Soluble Solid Content (SSC) in grapes using low-cost RGB sensors in challenging, uncontrolled field conditions. By creating a novel multi-season, multi-device dataset of Pizzutello Nero labeled for both SSC and color, we established a framework for developing and validating data-driven estimation models, addressing the practical limitations of expensive laboratory or hyperspectral equipment. While numerous studies explore laboratory equipment and hyperspectral data for SSC estimation through visual sensors, this study is among the few to focus exclusively on the practical scenario of RGB data collected directly in field conditions. The complexity of this task arises from covariate shifts inherent to the problem, including seasonal variations and differences in imaging devices.
Two algorithmic approaches were proposed: a lightweight method using classical handcrafted features, suitable for devices with limited computational power, and a multi-task deep neural network architecture that, while more computationally demanding, effectively leverages correlations between visual appearance (color and shades) and SSC.
Both methods were experimentally validated using field-collected data. Results indicate that the proposed techniques achieve reliable SSC estimation with a reasonable margin compared to the average human estimator. Our multi-task deep neural network achieved a best-performing Mean Absolute Error (MAE) as low as 1.05 °Brix on a challenging cross-device test set. This level of accuracy is highly competitive with reported results from in-field hyperspectral systems and is comparable to our human expert baseline, whose harvesting decisions the model could match with high precision and recall. The demonstrated robustness to seasonal and hardware-induced covariate shifts position our solutions as viable prototypes for low-cost, on-field applications.
@@
To foster further research we release our complete dataset, code, trained models, and training scripts used in this study. Future work will focus on extending the approaches to green cultivars, and on further studying how low-cost estimators can directly approximate hyperspectral devices in agronomic contexts..
CRediT authorship contribution statement
Thomas A. Ciarfuglia Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Investigation, Resources, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision, Project administration, Funding acquisition. Ionut M. Motoi: Software, Validation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization Leonardo Saraceni: Software, Validation, Data Curation Daniele Nardi: Funding acquisition, Writing - Review & Editing
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.